OpenStudy (anonymous):

Find the zeros of this quadratic. 2x^2 - 8 = 0

OpenStudy (anonymous):

\[2x ^{2}-8=0\]

OpenStudy (anonymous):

would it be x=4, 9

OpenStudy (anonymous):

@jim_thompson5910 sorry for troubling you so much

jimthompson5910 (jim_thompson5910):

2x^2 - 8 = 0 2x^2 = 8 x^2 = 8/2 x^2 = 4 x = ??? or x = ???

jimthompson5910 (jim_thompson5910):

also you can check your possible answers, checking x = 9 as a possible answer 2x^2 - 8 = 0 2(9)^2 - 8 = 0 ... plug in x = 9 2(81) - 8 = 0 162 - 8 = 0 154 = 0 ... which is FALSE So x = 9 is NOT a solution

OpenStudy (anonymous):

oops i meant to put x= 4, 0

jimthompson5910 (jim_thompson5910):

ok check x = 0 2x^2 - 8 = 0 2(0)^2 - 8 = 0 ... plug in x = 0 2(0) - 8 = 0 0 - 8 = 0 -8= 0 ... which is FALSE So x = 0 is NOT a solution

jimthompson5910 (jim_thompson5910):

see how I'm checking the answers?

OpenStudy (anonymous):

yeah thanks for all your help!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

wait, so how do i solve for the zeros. i got 4, 0 but those aren't true

jimthompson5910 (jim_thompson5910):

no I just showed you why it's not true I just showed how x = 0 wasn't a solution

jimthompson5910 (jim_thompson5910):

also I had steps leading up to this equation x^2 = 4 so if x^2 = 4, then x is equal to what two values?

jimthompson5910 (jim_thompson5910):

It's all above in separate posts

OpenStudy (anonymous):

ohh +- 2

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

to check each solution, just plug it in for each x and evaluate so you should be able to plug in x = 2 in for each x and evaluate and you will get zero if you do everything correctly that would confirm that x = 2 is a true solution to 2x^2 - 8 = 0

jimthompson5910 (jim_thompson5910):

I recommend doing this for any equation you solve because not all possible solutions are actual solutions